# MSViT: Improving Spiking Vision Transformer Using Multi-scale Attention Fusion

### v1

@@ Abstract
The combination of Spiking Neural Networks (SNNs) with Vision Transformer architectures has attracted significant attention due to the great potential for energy-efficient and high-performance computing paradigms. However, a substantial performance gap still exists between SNN-based and ANN-based transformer architectures. While existing methods propose spiking self-attention mechanisms that are successfully combined with SNNs, the overall architectures proposed by these methods suffer from a bottleneck in effectively extracting features from different image scales. In this paper, we address this issue and propose MSViT, a novel spike-driven Transformer architecture, which firstly uses multi-scale spiking attention (MSSA) to enrich the capability of spiking attention blocks. We validate our approach across various main data sets. The experimental results show that MSViT outperforms existing SNN-based models, positioning itself as a state-of-the-art solution among SNN-transformer architectures. The codes are available at https://github.com/Nanhu-AI-Lab/MSViT.
@@ 3 Methods
Figure 2: Comparison of three self-attention computation paradigms. (a) VSA employs floating-point matrix multiplication to assess the spatial correlation between Q and K, resulting in a computational complexity of $O(N^{2}D)$. (b) SSA lacks a dedicated temporal interaction module, maintaining the same complexity as VSA. (c) In contrast, MSSA introduces multi-scale interactions, maintaining the complexity to $O(T(ND)$).
@@ 4.2 Results on ImageNet-1K Classification
Comparing with ANN Models on ImageNet. MSViT is designed as an event-driven SNN model, in which the outputs of the embedding layers, the matrix calculation in the attention blocks, and information transmission are binary spikes $\{0,1\}$. As a result, the multiplications of the weight matrix and activations, such as RELUs, which are important, can be replaced by AND-ACcumulate (AC) operations, which benefit model training with high energy efficiency. Meanwhile, to achieve a competitive performance for ANN-based models, we apply the hierarchical transformer architecture to MSViT. Although the implementation of hierarchical architectures is relatively more complex than those using the same attention blocks throughout the networks (Zhou et al. (2023); Yao et al. (2024b)), it is still worth using a hierarchical architecture to reduce the performance gap between ANNs and SNNs. This is mainly because the hierarchical architecture naturally has the flexibility to model at various scales and has linear computational complexity with respect to image input size Liu et al. (2021). For instance, our MSViT outperforms the most well-known Transformer-based ANNs in performance with high energy efficiency under the same experiment conditions without pretraining or extra training data, among MSViT (69.80M, 85.06%, SNN, 45.88mJ), Swin Transformer (88M, 84.5%, ANN, 216.20mJ) Liu et al. (2021), DeiT-B (86M, 83.1%, ANN, 254.84mJ) Touvron et al. (2021) and ViT (85.59M, 77.9%, ANN, 254.84mJ) Dosovitskiy (2020) ( the results in detail are in Table 1).
@@ 6 Conclusion
In this work, we design a novel spike-driven multi-scale attention (MSSA), which involves a multi-scale feature fusion mechanism in the spiking hierarchical transformer architecture to improve the model’s performance. The multi-scale spiking attention fuses the low-level and high-level information, significantly improving the model performance with limited parameter increases. MSSA replaces the dot-product or Hadamard product existing in VSA or SSA with a column sum, maintaining the computation in linear complexity to the tokens of the inputs. Correspondingly, we also use Spiking Patch Embedding with Multi-scale Feature Fusion (SPEMSF) which enhances spiking representation for both high-level and low-level information, promoting improvement of the model. In addition, we customized a hierarchical spiking transformer architecture equipped with the aforementioned MMSA and SPEMSF. We name the model MSViT. Our model achieves state-of-the-art performance on both static and neuromorphic datasets. Notably, MSViT achieved top-1 accuracy over 85% with 69.80M parameters and $224^{2}$ by direct training. Leveraging the modelname’s superior capabilities, we strive to inspire confidence in the application of Spiking Neural Networks through our work.
Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2024YDLN0006), the Research Grants Council of the Hong Kong SAR (Grant No. PolyU25216423), and the Research Fund of the University of New South Wales (OP001-Z6666-PS75914).

### v3

@@ Abstract
The combination of Spiking Neural Networks (SNNs) with Vision Transformer architectures has garnered significant attention due to their potential for energy-efficient and high-performance computing paradigms. However, a substantial performance gap still exists between SNN-based and ANN-based transformer architectures. While existing methods propose spiking self-attention mechanisms that are successfully combined with SNNs, the overall architectures proposed by these methods suffer from a bottleneck in effectively extracting features from different image scales. In this paper, we address this issue and propose MSViT. This novel spike-driven Transformer architecture firstly uses multi-scale spiking attention (MSSA) to enhance the capabilities of spiking attention blocks. We validate our approach across various main data sets. The experimental results show that MSViT outperforms existing SNN-based models, positioning itself as a state-of-the-art solution among SNN-transformer architectures. The codes are available at https://github.com/Nanhu-AI-Lab/MSViT.
@@ 3 Methods
Figure 2: Comparison of three self-attention computation paradigms. (a) VSA employs floating-point matrix multiplication to assess the spatial correlation between Q and K, resulting in a computational complexity of $O(N^{2}D)$, (b) SSA lacks a dedicated temporal interaction module, maintaining the same complexity as VSA and, (c) MSSA introduces multi-scale interactions, reducing the complexity to $O(T(ND)$).
@@ 4.2 Results on ImageNet-1K Classification
Comparing with ANN Models on ImageNet. MSViT is designed as an event-driven SNN model, in which the outputs of the embedding layers, the matrix calculation in the attention blocks, and information transmission are binary spikes $\{0,1\}$. As a result, the multiplications of the weight matrix and activations, such as RELUs, which are important, can be replaced by AND-ACcumulate (AC) operations, which benefit model training with high energy efficiency. Meanwhile, to achieve a competitive performance for ANN-based models, we apply the hierarchical transformer architecture to MSViT. Although the implementation of hierarchical architectures is relatively more complex than those using the same attention blocks throughout the networks (Zhou et al. (2023); Yao et al. (2024b)), it is still worth using a hierarchical architecture to reduce the performance gap between ANNs and SNNs. This is mainly because the hierarchical architecture naturally has the flexibility to model at various scales and has linear computational complexity with respect to image input size Liu et al. (2021). For instance, our MSViT outperforms the most well-known Transformer-based ANNs in performance with high energy efficiency under the same experiment conditions without pretraining or extra training data, among MSViT (69.80M, 85.06%, SNN, 45.88mJ), Swin Transformer (88M, 84.5%, ANN, 216.20mJ) Liu et al. (2021), DeiT-B (86M, 83.1%, ANN, 254.84mJ) Touvron et al. (2021) and ViT (85.59M, 77.9%, ANN, 254.84mJ) Dosovitskiy (2020) ( the detailed results are in Table 1).
@@ 6 Conclusion
In this work, we design a novel spike-driven multi-scale attention (MSSA), which involves a multi-scale feature fusion mechanism in the spiking hierarchical transformer architecture to improve the model’s performance. MSSA fuses low-level and high-level information of inputs, significantly improving the model performance with limited parameter increases. Furthermore, MSSA replaces the dot-product or Hadamard product existing in VSA or SSA with a column sum, maintaining the computation in linear complexity to the tokens of the inputs. Correspondingly, we also utilize Spiking Patch Embedding with Multi-scale Feature Fusion (SPEMSF), which enhances spiking representation for both high-level and low-level information of inputs, thereby promoting model improvement. Finally, we implement a hierarchical spiking transformer architecture equipped with the aforementioned MMSA and SPEMSF, namely MSViT. We have conducted extensive experiments, and the results show that our model achieves state-of-the-art performance on both static and neuromorphic datasets. Notably, MSViT achieved top-1 accuracy on ImageNet-1K over 85% with 69.80M parameters and image input of size $224^{2}$ by direct training from scratch. Leveraging the MSViT’s superior capabilities, we strive to inspire confidence in the application of Spiking Neural Networks through our work.